OpenStudy (anonymous):

66 2/3% as a fraction

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2%2F3

OpenStudy (anonymous):

2/3=0.666666666666666666 So 66 and 2/3=66.666666666666666666 66.666666666666666666666666.../100 (percent= per-100)=0.666666666666666666666666...=2/3